# Liwen Wang1,2,  Yuanyuan Yuan2,  Ao Sun2,  Zongjie Li2,  Pingchuan Ma2,  Daoyuan Wu2,  Shuai Wang2

### v2

@@ 7.1 Alignment and Faithfulness Evaluation
Table 7 presents the results of alignment and faithfulness evaluations for four MLLM-enabled mutations. Different from our human studies, the two numerical metrics cannot distinguish the alignment and faithfulness of different MLLMs, despite that they are widely adopted to benchmark MLLMs [53, 68, 28]. The cosine similarity values in Table 7 range from -1 to 1, and the CLIP embedding often cannot capture relations between objects in the text [63, 70]. In that sense, the moderately high alignment results (all are positive and are greater than 0.4) in Table 7 indicate that these “semantic-additive” mutations do generate the new semantics. However, since MLLMs have varying capabilities to comprehend relations among objects, they may put the added semantics into a wrong position. For example, as in Fig. 3, when implementing Legs, GPT-4V makes the person’s legs bionic rather than the dog, which can be identified by humans and results in lower alignment scores in Table 3. All MLLMs have almost perfect faithfulness results in Table 7, indicating that MLLMs are capable of limiting their implemented mutations within the mutated subjects or background semantics.
Table 8: Faithfulness results of Eyes.
LLaVA CogVLM GPT-4V InstP2P Traditional
MSE 17.97 85.27 18.03 1014.97 200.79
Perceptual-level Mutation (⑥). Perceptual-level mutations also lack ground truth and reference images. Although the anticipated mutation result of Eyes can be described as “a person with closed eyes”, we find that the CLIP embeddings mentioned above cannot capture such detailed eye status in both text and image. Therefore, we only evaluate the faithfulness for Eyes. Because eyes in FFHQ dataset locate in a unified region, we can easily mask out the eyes and check if the remaining pixels are modified. The pixel difference is calculated using MSE; a lower value indicates a better faithfulness.
Interestingly, Table 8 seemingly contradicts our human study results: the MSE of LLaVA and GPT-4V are much lower than traditional method, whereas human participants rate traditional method as the best in the faithfulness evaluation. Consider the mutation examples in Fig. 3, LLaVA- and GPT-4V-implemented mutations indeed replace eyes in the seed image, which limits the mutated region within the eyes but produces unnatural images. Such unnaturalness negatively impacts the faithfulness under human’s view. In contrast, the traditional method slightly edits other relevant semantics (e.g., muscles around the eyes) to generate naturally mutated images. Despite that pixels outside the eyes are also modified, these changed pixels are unaware under human observation.
7.2 Validity Evaluation
Since MLLMs cannot implement pixel-, geometrical-, and perceptual-level mutations as evaluated in Sec. 6.2 and Sec. 7.1, we use numerical metrics to evaluate the validity for style-level (③-⑤) and MLLM-enabled mutations (⑦-⑩).
Table 9: Validity results of style-level mutations.
IS FID Precision Recall
LLaVA 18.26 120.49 0.33 0.29
CogVLM 17.11 124.30 0.34 0.34
GPT-4V 17.61 123.70 0.39 0.31
InstP2P 20.17 127.95 0.35 0.24
Traditional 19.84 124.51 0.46 0.19
Table 10: Validity results of MLLM-enabled mutations.
IS FID Precision Recall
LLaVA 14.51 88.09 0.58 0.33
CogVLM 13.71 89.65 0.60 0.34
GPT-4V 14.82 88.67 0.53 0.36
InstP2P 23.93 93.95 0.58 0.35
IS and FID Metrics. Following existing works, we take two popular metrics, Inception Score (IS) [51] and Frechet Inception Distance (FID) [29], to measure the validity of the mutated images. Both IS and FID use a set of real images as the reference, and compare the mutated images with these reference images to estimate how likely the mutated images are real. A lower FID value indicates a better realism, whereas a higher IS value indicates a better realism.
Precision & Recall. Although IS/FID are widely used by image generation and VDL testing to assess how realistic an image is [26, 65], they jointly consider the validity and diversity. Some of our evaluated mutations are non-deterministic (e.g., style-level mutations), where multiple (possible) mutated images can be produced using one seed image. Thus, a better IS or FID result may be due to diversified mutations, not a better validity. We use the Precision & Recall metrics [50, 39] to separately evaluate the validity and diversity, respectively. The Precision & Recall here are different from the conventional precision/recall that count positive and negative results; they are new metrics proposed to assess image quality. Both Precision & Recall range within $[0,1]$ and higher is better.
Table 9 shows results of the validity evaluation for style-level mutations implemented using traditional methods and different MLLMs. Considering the high magnitude of FID values, both IS and FID results cannot notably distinguish different methods. When separately evaluating the validity and diversity, traditional method has the highest validity (Precision) but the lowest diversity (Recall). That is, MLLMs are superior in diversifying the seed images with style-level mutations, whereas traditional method can better preserve the validity.
Validity results of MLLM-enabled mutations are given in Table 10. Only IS values can distinguish different MLLMs and InstP2P has the best IS result, which is consistent with our human study results where InstP2P largely outperforms other MLLMs in the validity evaluation of Clothes and Legs. By corss-comparing results in Table 10 and Table 9, especially FID, Precision, and Recall values, we can see that MLLMs perform better on the semantic-additive mutations than on the style-level mutations, as supported by both better validity and diversity (note that lower FID is better).
[…]
VDL Model Contrast Rotation Artifying SnowG Background Clothes Legs Glasses
ConvNeXt LLaVA N/A N/A 7.6% 4.8% 5.2% 7.7% 2.8% 7.7%
CogVLM 9.6% 4.8% 7.6% 10.1% 4.5% 8.6%
GPT-4V 9.2% 3.6% 6.8% 9.7% 1.6% 5.7%
InstP2P 6.4% 4.4% 11.2% 1.2% 1.2% 0.8%
Traditional 6.0% 0.4% 6.8% 14.4% N/A
Avg. 6.0% 0.4% 7.9% 6.4% 7.7% 7.2% 2.5% 5.7%
SwinTransformer LLaVA N/A N/A 8.0% 4.8% 5.2% 7.3% 2.8% 7.7%
CogVLM 9.6% 4.8% 7.2% 10.9% 4.5% 9.4%
GPT-4V 9.2% 4.4% 7.2% 9.3% 1.6% 6.1%
InstP2P 5.6% 4.0% 13.2% 2.8% 1.6% 2.0%
Traditional 6.8% 0.4% 8.4% 16.0% N/A
Avg. 6.8% 0.4% 8.2% 6.8% 8.2% 7.6% 2.6% 6.3%
ViT LLaVA N/A N/A 8.0% 4.8% 4.4% 7.7% 3.2% 6.1%
CogVLM 9.6% 5.2% 7.6% 10.5% 4.9% 9.8%
GPT-4V 7.6% 4.4% 7.6% 10.1% 0.8% 5.3%
InstP2P 7.2% 3.6% 12.0% 2.0% 1.6% 1.6%
Traditional 8.4% 0.4% 7.2% 16.0% N/A
Avg. 8.4% 0.4% 7.9% 6.8% 7.9% 7.6% 2.6% 5.7%
7.3 Testing Effectiveness Evaluation
We also employ mutated images generated from different mutations to test VDL models and evaluate their effectiveness in triggering VDL faults. Given the wide adoption of VDL systems, it is impractical to consider all VDL applications. We therefore choose different VDL-based classifiers as one representative application.
We choose three recent SoTA VDL-based classifiers, Vision Transformer (ViT) [17], SwinTransformer (SwinT) [44], and ConvNeXt [45]. These models are officially released by PyTorch and are trained with ImageNet, where SnowD (⑤) and Eyes (⑥) are inapplicable. Mutated images are fed into a popular VDL testing tool DeepHunter [62] to detect VDL faults. DeepHunter implements a checker to rule out potentially broken, invalid mutated images to reduce false positive VDL faults. We configure its checker such that invalid images can be mostly filtered out, per our observation.
Results are given in Table 11. While the same seed images are used, images generated using different mutations have distinct effectiveness when triggering VDL faults. Pixel-level (①) and style-level mutation (③ and ④) have high effectiveness on average. This is consistent with the texture bias revealed in [24]: modern image classifiers are sensitive to perturbations of texture in images. Here, texture refers to the appearance of an object’s surface, such as the intensity and the arrangement of colors. Both pixel- and style-level mutations modify texture-related semantics and can more frequently trigger VDL faults.
Interestingly, while SnowG (④) implemented using traditional method mostly does not mutate the seed image (see Sec. 6.2 and Sec. 7.1), their unperceivable perturbations can effectively mislead the tested VDL model. Despite the high effectiveness, considering that SnowG is not accurately implemented, these mutated images should be less valuable as they give a false sense that the tested VDL model is sensitive to weather conditions. For the four MLLM-enabled mutations (⑦-⑩), Background (⑦) and Clothes (⑧) have the highest effectiveness. Given that Background does not alter the subject based on our results in Sec. 6.2 and Sec. 7.1, we suspect that these SoTA VDL models may largely rely on background semantics to recognize the subject. Clothes often generates unusual combinations of semantics, e.g., a dog wearing clothes may less likely to appear in these models’ training data. Thus, it is reasonable that Clothes is more effective to challenge the tested VDL models. Legs (⑨) and Glasses (⑩) are less effective because their mutated semantics (i.e., legs and eyes) are not crucial to recognize dogs.
8 Lessons, Discussion, and Threat To Validity
@@
New Mutations. Our evaluation shows that MLLMs are suitable for mutations transferring semantics between images, e.g., style-level mutations that transfer styles, and semantic-additive mutations that add new semantics (from other images) to the seed. These are highly valuable in VDL testing, as they can generate diverse and valid test inputs, and importantly, these mutations are hardly feasible with traditional methods. We thus believe it is high time to incorporate MLLMs into VDL testing tasks, which shall enable comprehensive and finer-grained testing of VDL systems where subtle semantic changes are critical and can lead to catastrophic failures, such as in autonomous driving and drone navigation. Nevertheless, contemporary MLLMs are less capable of editing existing semantics. That said, despite the optimism, MLLMs are not a panacea for all VDL testing tasks. We thus advocate to integrate and carefully choose MLLMs and traditional methods in VDL testing.
Different MLLM Pipelines. By comparing MLLM pipelines (as introduced in Sec. 5.2) and considering semantic-additive mutations, the unidirectional and fine-tuning-based pipeline appears more applicable to silent mutations that modify the image background or subtly alter the image subject. The bidirectional pipeline is superior in actively editing the subject. These findings suggest that different MLLM pipelines have distinct capabilities in reasoning the relations among objects, leading to varied performance for different mutations. Therefore, we suggest users choose the MLLM pipeline based on the mutation’s extent and the testing target (e.g., a VDL’s robustness to varied backgrounds or subjects). Moreover, this finding highlights the importance of understanding and decomposing scenes in seed images before mutating subjects or backgrounds; emerging research sheds light on this direction [54, 69].
Threat to Validity. The human study may be biased, and to mitigate this threat, we have involved 20 participants with relevant expertise. We presume that it is sufficient to guarantee the credibility of our study. We also conducted a quantitative evaluation to complement the human study. Second, the numerical criteria used in the quantitative evaluation may not fully capture the quality of the generated images. To mitigate this, we follow the existing criteria (e.g., FID and IS) used in the literature. Moreover, our findings may not be generalizable; we mitigate this by using diverse sets of images, mutation methods, MLLMs, and VDL systems.
@@ 9 Conclusion
This paper presents an in-depth study on MLLM-produced test inputs for VDL testing. We categorize image mutation methods and evaluate their quality using several criteria. We provide insights into the (in-)adequacy of MLLMs for VDL testing and advocate for the careful integration of MLLMs and traditional image mutation methods for better testing effectiveness and application scenarios.

### v3

@@ 7.3.2 Embedding-Based Metrics
Table VII presents the results of alignment and faithfulness evaluations for four MLLM-enabled mutations using popular embedding-based metrics. The alignment results of Background are relatively better than that of Clothes, Legs, and Glasses, indicating that the metric can detect the mutated background, but is unable to detect the mutated finer-grained semantics on subjects. The high faithfulness results show that the embedding-based metric can reflect the (ought-to-be) unchanged semantics. Nevertheless, unlike our human studies that show notable differences in alignment and faithfulness results of different MLLMs, the two numerical metrics cannot distinguish their varied results, despite that similar methods show promising results for traditional mutations [51].
Note that traditional mutations like style- and perceptual-level mutations only apply to specific domains, and accordingly, their corresponding image embeddings can be limited to the target domain to improve precision. E.g., when measuring mutations applied to face photos, existing works leverage face generators to obtain face embeddings and do not consider non-face images. Similarly, when measuring the achieved style-level mutations, prior works leverage style extractor to specifically compute the style distance. However, MLLM-enabled mutations are applicable to more general images. Although CLIP can obtain embeddings for diverse images from different domains, these embeddings consequently sacrifice their precision; similar observations are also noted in recent research [79, 80].
Answer to RQ3: Prior validation metrics are less applicable to MLLM-enabled mutations. Validity metrics like IS and FID jointly consider validity and diversity, whose numerical results can be misled by more diversified test inputs generated via MLLMs. Adapting metrics that separate validity and diversity (e.g., the Precision & Recall metric [74, 75]) for MLLMs may be a good starting point. Embedding-based metrics also become inaccurate due to the broader application scopes of MLLM-enabled mutations. Considering the high challenge of designing unified and accurate embedding models for diverse images, leveraging MLLMs to generate textual descriptions for mutants (following the bidirectional pipeline in Sec. 5.2) may be promising to measure the achieved extents of MLLM-enabled mutations.
8 Discussion and Threat To Validity
@@
New Mutations. Our evaluation shows that MLLMs can generalize mutations transferring semantics between images, e.g., style-level mutations that transfer styles, and enable semantic-replacement mutations that add new semantics (from other images) to the seed. These are highly valuable in VDL testing, as they can generate diverse and valid test inputs, and importantly, these mutations are largely limited or hardly feasible with traditional methods. We thus believe it is high time to incorporate MLLMs into VDL testing tasks, which shall enable comprehensive and finer-grained testing of VDL systems where subtle semantic changes are critical and can lead to catastrophic failures, such as in autonomous driving and drone navigation.
Nevertheless, contemporary MLLMs are less capable of editing existing semantics. That said, despite the optimism, MLLMs are not a panacea for all VDL testing tasks. We thus advocate to integrate and carefully choose MLLMs and traditional methods in VDL testing.
Threat to Validity. The human study may be biased, and to mitigate this threat, we have involved 20 participants with relevant expertise, and each question is evaluated by different participants to reduce personal bias. We also conduct a sanity check to rule out answers from participants who likely misunderstand the questions. We presume that it is sufficient to guarantee the credibility of our study.
Second, we also study the specialties of different pipelines of employing MLLMs for mutations. All two available pipelines are considered in our evaluations, and we incorporate four SoTA MLLMs into them to implement 10 representative mutations. We believe the evaluated cases are comprehensive to alleviate potential bias.
Third, the suitability of existing validation metrics for MLLM-based mutations are also evaluated. To make the selected metrics representative, we consider metrics that are widely adopted and have been proven effective in existing testing works. We believe this setup is sufficient to reflect the common issues in validation metrics tailored for traditional mutations.
@@ 9 Conclusion
This paper presents an in-depth study on MLLM-based mutations in VDL testing and reveals the complementary positioning of MLLMs and traditional mutations. We provide insights into the (in-)adequacy of MLLMs for VDL testing, the specialties of different mutation pipelines of MLLMs, and the limitations of existing validation metrics when facing MLLM-based mutations. Our findings call for careful integration of MLLMs and traditional image mutations for better testing effectiveness and application scenarios.